# Impact of projection-induced optical selection bias on the weak lensing mass calibration of galaxy clusters

### v1

@@ V Summary and discussion
Our findings indicate that the projection-induced selection bias is the most detrimental among the known systematic biases impacting cluster mass calibration. Other known systematics include the impact of baryons on the mass distribution [67, 68, 69, 70, 71, 72, 73, 74, 75], the offset between the inferred cluster centers and the true halo centers (miscentering) [76, 77, 78, 79], and the diluted lensing signal due to cluster members (the so-called boost factor) [77, 65]. Ref. [80] has assessed the impact of these systematics on the weak lensing mass bias for $\sim 1000$ clusters selected in mm-wave by the South Pole Telescope, which is unaffected by the optical selection bias discussed in this paper. Their weak lensing mass bias is dominated by baryonic effects (for $z\lesssim 0.4$) and the photometric redshift calibration of source galaxies (for $z\gtrsim 0.4$); see their Table IV and Fig. 10. These systematics lead to a mass bias around $5-10\%$; their highest bias is 16%, which occurs at $z\approx 1$. These systematics also impact optically selected clusters, but they are subdominant compared with the optical selection bias discussed in this paper.
How do we constrain the projection-induced selection bias using observational data? Spectroscopic redshifts of member galaxies can help us separate projected vs. physically-associated cluster members [30, 33]. Cross-correlating clusters with spectroscopic galaxies can help us calibrate the redshift distribution of cluster members [73, 81]. Ref. [29] also shows how to use photometric galaxies to calibrate projection effects. Combining cluster lensing and cluster–galaxy cross correlations can help us self-calibrate the selection bias, which boosts all cluster-related 2-point statistics [82, 83, 17, 42]. Multiwavelength analyses comparing cluster lensing signals of optically selected clusters and gas-selected clusters (observed in mm-wave or x-ray) can help us calibrate the lensing bias [84]. In sum, we advocate a multiprobe, multiwavelength approach for cluster cosmology. Future modeling strategies should focus on self-consistently modeling the impact of projection on richness, lensing, and gas observables [25, 85, 86].
How do we treat the selection bias in optical cluster cosmology analyses? Recent studies have modeled the selection bias as functions of radius, richness, and redshift [14, 17, 87, 88]. Another promising route is to use $N$-body simulations to forward-model richness and lensing, an approach that automatically takes into account the selection bias [43, 31, 89]. Since the mass bias depends on the galaxy model, the cluster finding, and the radial range used, we will need a flexible model, which can be achieved with simulation-based forward modeling. We anticipate that the simulation-based forward modeling has the potential to help us realize the constraining power of optical clusters.

### v2

@@ V Summary and discussion
Our findings confirm that the projection-induced selection bias is the most detrimental of the known systematic biases impacting cluster mass calibration. Other known systematics include the impact of baryons on the mass distribution [Schaller15, Henson17, Lee18, Cromer21, GiriSchneider21, Grandis21b, Sunayama23, To24baryon, Dalal25], the offset between the inferred cluster centers and the true halo centers (miscentering) [Johnston07, Kohlinger15, 37, Sommer22], and the diluted lensing signal due to cluster members (the so-called boost factor) [Kohlinger15, 32]. Ref. [Bocquet24a] has assessed the impact of these systematics on the weak lensing mass bias for $\sim 1000$ clusters selected in millimeter-wave by the South Pole Telescope, which is unaffected by the optical selection bias discussed in this paper. Their weak lensing mass bias is dominated by baryonic effects (for $z\lesssim 0.4$) and the photometric redshift calibration of source galaxies (for $z\gtrsim 0.4$); see their Table IV and Fig. 10. These systematics lead to a mass bias around $5-10\%$; their highest bias is 16%, which occurs at $z\approx 1$. These systematics also impact optically selected clusters, but they are subdominant compared with the optical selection bias discussed in this paper.
How do we constrain the projection-induced selection bias using observational data? Spectroscopic redshifts of member galaxies can help us separate projected vs. physically-associated cluster members [18, Myles25projection]. Cross-correlating clusters with spectroscopic galaxies can help us calibrate the redshift distribution of cluster members [Sunayama23, Yang25]. Ref. [9] also shows how to use photometric galaxies to calibrate projection effects. Combining cluster lensing and cluster–galaxy cross correlations can help us self-calibrate the selection bias, which boosts all cluster-related 2-point statistics [To21a, 29, Sunayama23, 36]. Multiwavelength analyses comparing cluster lensing signals of optically selected clusters and gas-selected clusters (observed in mm-wave or x-ray) can help us calibrate the lensing bias [Zhou24]. In sum, we advocate a multiprobe, multiwavelength approach for cluster cosmology. Future modeling strategies should focus on self-consistently modeling the impact of projection on richness, lensing, and gas observables [2, Shirasaki16, Ragagnin25].
How do we treat the selection bias in optical cluster cosmology analyses? Recent studies have modeled the selection bias as functions of radius, richness, and redshift [29, Sunayama23, DESY3CL, To25]. Another promising route is to use $N$-body simulations to forward-model richness and lensing, an approach that automatically takes into account the selection bias [24, 14, Cao26unify]. Since the mass bias depends on the galaxy model, the cluster finding, and the radial range used, we will need a flexible model, which can be achieved with simulation-based forward modeling. We anticipate that the simulation-based forward modeling has the potential to help us realize the constraining power of optical clusters.